# An X-ray absorption spectrum database for iron-containing proteins

### v1

@@ Methods
In addition to raw data comparison, we assessed the consistency of the annotations provided by the experts using the Intra-class Correlation Coefficient (ICC)[22, 23]. This analysis is crucial for understanding the reliability of the annotations across different experts.
  • Annotation Work. The same raw data-plot dataset was used for the initial annotation work by the four experts. Using an identical dataset ensures that the comparison and consistency evaluations are based on identical data points, making the validation more coherent and reliable.
  • ICC Calculation: Based on the annotation work, ICC was calculated for each data point in the sequences and averaged to provide an overall measure of agreement.
Table 4 presents results from an Intraclass Correlation Coefficient (ICC) analysis, which assesses the reliability of measurements made by multiple raters. The ICC types include ICC1 (Single rater’s absolute agreement) which evaluates the consistency of measurements from each rater as if they are the only one measuring; ICC2 (Single random raters’ agreement) which assumes that the raters are randomly selected from a broader population, thus the findings are generalized beyond the specific raters used; and ICC3 (Single fixed rater’s agreement) that treats the raters as the only ones of interest without intention to generalize to others, thus accounting for systematic differences among them. The ’k’ variants (ICC1k, ICC2k, ICC3k) improve reliability by using the average of measurements across raters. All ICC values are larger than 0.91, and some of them approach 1, indicating superior reliability. The substantial F-statistics and near-zero p-values provide strong statistical evidence against the null hypothesis of no rater consistency. Degrees of freedom, df1 and df2, relate to the numerator and denominator of the F-test, respectively. The 95% confidence intervals, narrowly focused around 0.95, underscore the measurement precision. This comprehensive evaluation is pivotal in domains demanding high reliability, such as health sciences and psychological assessments, ensuring consistent and replicable measurements across different studies and settings.
Documentation and Transparency:
Another spreadsheet records the processes involved in extracting iron XAS spectra. It highlights various challenges encountered, such as missing axes on figures and ambiguities in figure interpretation. Such detailed records are crucial for enabling future researchers to understand and evaluate the decision-making processes affecting data inclusion or exclusion. Notebly, we includes and distinguishes between EXAFS data represented in both k-space and r-space. This enhancement ensures that researchers can access and utilize the specific type of data they require for their analyses. We have meticulously labeled each dataset entry in our documentation files to reflect this categorization, addressing a critical need for clarity that was missing in earlier versions of our paper. By doing so, we have significantly improved the utility and accuracy of our dataset, facilitating more precise and informed scientific research in the field of X-ray Absorption Spectroscopy. The details of documentation information are explained in Table 6, besides, Figure 5(b) shows an example of XAS spectrum data documentation, revealing the key information and issues recorded during the annotation work.
Data entries within our datasets were cross-verified against original publications to ensure accuracy and reliability. Spectral and structural data were, wherever feasible, compared against established databases to ensure consistency. Discrepancies and significant findings are carefully documented. The methodologies employed for data extraction are thoroughly detailed, including the tools used, such as the WebPlotDigitizer[20] for converting plot images to numerical data. Each dataset entry is accompanied by extensive metadata such as detailed URLs of original publications, covering all aspects from data extraction to final dataset compilation. This approach not only enhances reproducibility but also facilitates the integration of our data with other datasets. It ensures a clear understanding of the context and specifics of the data collected.
Usage Notes
Software Packages and Tools
For analyzing the X-ray Absorption Spectroscopy (XAS) data provided in this dataset, we recommend using ATHENA[24] from the Demeter software suite, which is widely used for processing XAS data. Additionally, the WebPlotDigitizer[20] tool (version 4.2) was used to extract numerical data points from spectral plots, and we suggest researchers use this tool if they need to replicate the data extraction process.
Data Normalization and Processing
The XAS spectra in this dataset have been standardized through interpolation to ensure lengths—100 data points for XANES and EXAFS. Before performing further analysis or comparison with other datasets, appropriate normalization techniques, such as edge-step normalization should be applied. For advanced processing, employing machine learning techniques to explore relationships between spectral features and protein local structures is recommended.
Integration with Other Datasets
This dataset is designed to integrate seamlessly with other XAS databases like XASLIB[11] or the Materials Project[12]. The focus on iron-containing proteins ensures high specificity, which can be directly compared with similar spectra from inorganic compounds in other databases.
Reuse and Reproducibility
We encourage the reuse of this dataset in machine-learning applications, such as predicting protein local structure from XAS spectra or vice versa. All data records are structured to support easy import into computations. To ensure reproducibility, we suggest maintaining comprehensive documentation when applying this dataset to further applications, especially when combining this dataset with other data sources.
Code availability
@@
The code used for processing the spectra and molecular structures is available at https://github.com/Airscker/XDIP. Including scripts for data interpolation, normalization, and basic spectral analysis, providing a starting point for researchers who are interested in further processing the data. The software used for extracting numerical points from spectrum plots, i.e., WebPlotDigitizer V4, is open-source and available at https://github.com/automeris-io/WebPlotDigitizer.
Acknowledgements
@@
The work was supported in part by the Physical Biosciences Program and the Photochemistry and Biochemistry group within the US Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences (KC030402).
Author contributions statement
All authors contributed substantively to the work presented in this paper. Yufeng Wang led the research efforts, including XAS extraction, literature retrieval, data processing, and manuscript writing. Peiyao Wang contributed to literature retrieval, XAS data annotation, and data cleaning. Emerita Mendoza Rengifo and Dali Yang performed protein structure annotation and provided significant input in manuscript preparation. Keying Jia and Rutwik Segireddy assisted with XAS extraction. Lu Ma, Yuewei Lin, Qun Liu, and Haibin Ling provided mentorship throughout the study, with Qun Liu and Haibin Ling serving as corresponding authors.
Competing interests
@@
Figure 1: The schematic overview of dataset construction pipeline. (a) The literature search process, where keywords constructed from two sets (shown at the top of the panel) were used to search related literature across various paper databases. (b) The data processing workflow conducted by human experts. From the literature, we first identified Fe-related protein structures and their corresponding absorption spectra, converting this text/plot information into computer-readable and processable data. Subsequently, the dataset was refined by removing samples with issues such as missing structures or spectra. (c) An overview of the extracted database, where each sample includes the protein’s (local) structure, its corresponding absorption spectrum of the Fe element, and the basic information of the paper from which the data were derived.
Figure 2: Distribution of XANES and EXAFS Spectrum Lengths. The histogram displays the distribution of spectrum lengths for XANES (green) and EXAFS (orange) datasets.
Figure 3: Structure of data sample. For each paper, there may be multiple extractable information, so we stacked the spectrums and their corresponding structures. (a) The literature metadata, which includes every paper’s DOI, title, and absorbing element. In this study, we focus exclusively on the Fe element. (b) The extracted spectrum data. We categorized the absorption spectrums into two categories: near-edge absorption spectrum and extended absorption spectrum.(c) The structure extracted from the paper. Sometimes, the structure of the target material is available online by searching its ID or name, or it can be directly accessed from its SMILES[25, 26, 27] representation. Otherwise, we constructed the protein structure by labeling its adjacent matrix, atom list, atom coordinates, bond lengths, bond angles, and optional notes.
(a) Distribution of annotation MSE for XANES samples across experts
(b) Distribution of annotation MSE for EXAFS samples across experts
Figure 4: Distribution of annotation MSE for XAS samples across experts.
(a) Example of protein structure data documentation
(b) Example of XAS spectrum data documentation
Figure 5: Examples of protein structure and XAS spectrum data documentation. (a) This example presents a data entry from our database, including the paper’s title, DOI, and URL. To accurately document the data, we included sections for both XANES and EXAFS to help identify the spectrum type corresponding to the protein structure. Due to missing descriptions in some papers, certain annotation elements, such as standards (which refer to the standard sample used) and ligation details (which provide information on ligands or binding groups associated with the protein structures), are omitted. Despite the missing information, we listed the samples mentioned in the paper that have both spectra and structures, with these key points shown in the sample section under the XANES and EXAFS parts. As outlined in the Data Records section, fields such as SMILES, PDB, and CCDC are selectively filled. (b) This table provides an example of an XAS spectrum data annotation log. We recorded the title of the literature, the location of the iron spectrum plot, its DOI, and URL. In addition to these key details, we included comments on the quality or issues identified in the spectral or structural data, such as missing axes or unclear lines. Furthermore, to ensure the accuracy and usability of our data, we highlighted the type of EXAFS spectrum extracted from the literature, noting whether the article presented Fourier-transformed EXAFS or not.
@@
Table 1: Number of papers searched by using different keyword combinations
@@
The matrix of bond lengths between chemical bonded atoms Bond Lengths matrix of float numbers
The list of bond angles among neighborhood bonds Bond Angles dictionary whose key is the angle vertex and the value is float numbers
@@
Table 2: Format of each data record: description, key label, mandatory, data type.
Expert 1 Expert 2 Expert 3 Expert 4 Average
XANES MSE 0.0161 0.0014 0.0007 0.0521 0.0176
EXAFS MSE 0.0946 0.0928 0.1015 0.0944 0.0958
Table 3: MSE values of each expert.
Type Description ICC F df1 df2 pval CI95%
ICC1 Single raters absolute 0.91382703 15414.6102 99 300 1.39E-06 [0.898,0.931]
ICC2 Single random raters 0.91948134 21277.5769 99 297 5.57E-14 [0.876,0.947]
ICC3 Single fixed raters 0.94636021 21277.5769 99 297 5.57E-14 [0.933,0.959]
ICC1k Average raters absolute 0.96128244 15414.6102 99 300 1.39E-06 [0.948,0.972]
ICC2k Average random raters 0.96611623 21277.5769 99 297 5.57E-14 [0.927,0.982]
ICC3k Average fixed raters 0.98083997 21277.5769 99 297 5.57E-14 [0.974,0.986]
Table 4: ICC values for inter-expert annotation work.
@@
CCDC
Cambridge Crystallographic Data Centre number, providing a reference to
crystallographic data related to the protein.
Table 5: Description of protein database documentation elements
@@
Comments
Notes or comments about the quality or issues identified in the
spectral or structural data, such as missing axes or unclear lines.
EXAFS Type
Specifies the type of EXAFS data presented,
whether it is raw signal data or Fourier transformed data.

### v2

@@ 4 Technical Validation
To further evaluate annotation consistency, we computed the Intra-class Correlation Coefficient (ICC)koch2004intraclass ; koo2016guideline , a widely used metric for inter-rater reliability. The process involved:
  • Annotation Dataset. The same synthetic dataset was independently annotated by all four experts, ensuring that all raters worked on identical data.
  • ICC Calculation. We computed ICC values for each point in the interpolated spectra and averaged them to obtain overall agreement metrics across expert annotations.
Table 2: Intraclass Correlation Coefficient (ICC) values for inter-expert agreement.
Type Description ICC F df1 df2 95% CI
ICC1 Single rater (absolute) 0.914 15414 99 300 [0.898, 0.931]
ICC2 Single rater (random effects) 0.919 21277 99 297 [0.876, 0.947]
ICC3 Single rater (fixed effects) 0.946 21277 99 297 [0.933, 0.959]
ICC1k Average of raters (absolute) 0.961 15414 99 300 [0.948, 0.972]
ICC2k Average (random effects) 0.966 21277 99 297 [0.927, 0.982]
ICC3k Average (fixed effects) 0.981 21277 99 297 [0.974, 0.986]
Table 2 summarizes the ICC results. All values exceed 0.91, with several above 0.96, indicating excellent inter-expert reliability. The strong F-statistics and tight 95% confidence intervals confirm that the observed agreement is statistically significant and not due to chance. This level of consistency is particularly important in scientific applications like XAS interpretation, where subtle deviations can lead to different conclusions about structural or electronic properties.
Data Integrity Through Documentation and Transparency
To ensure reproducibility and traceability, we implemented a comprehensive documentation protocol. Each data record is accompanied by detailed metadata, including extraction notes, structural identifiers (e.g., PDB IDs), and reference sources.
Ambiguities encountered during data extraction—such as missing axes or unclear line styles—were explicitly recorded. We also documented whether EXAFS spectra were presented in $k$-space or $r$-space, which can significantly affect downstream analysis. All entries were cross-validated against their source publications and, where applicable, public databases.
This rigorous documentation pipeline not only ensures data quality and transparency but also enables seamless integration with other datasets. Further examples of our documentation format are provided in Supplementary Section A.4.
5 Baseline Models
To demonstrate the utility of our curated XAS dataset and establish a performance benchmark for future research, we evaluate two complementary predictive tasks: (1) predicting Fe K-edge XAS spectra from local atomic structures, and (2) inferring key structural properties from the spectra. This dual approach validates the dataset’s richness for both forward and inverse modeling problems.
5.1 Task 1: Structure-to-Spectrum Prediction
The first task learns a mapping function $f:G\rightarrow S$, where $G=(V,E)$ is a graph representing the local atomic environment of the iron center, and $S$ is the corresponding XAS spectrum. The local structure of each iron center is converted into an input graph, where atoms are nodes ($v\in V$) and chemical bonds are edges ($e\in E$). Node features include atomic properties like element type and mass, while edge features can represent bond types. The output target, $S$, is the standardized XAS spectrum (both XANES and EXAFS) interpolated to a uniform length of 100 data points.
For this task, we employ three widely-used Graph Neural Network (GNN) architectures, each comprising several graph convolutional layers, a graph pooling layer, and a multi-layer perceptron (MLP) head that maps the final graph representation to the 100-point spectrum vector. The models are: a foundational Graph Convolutional Network (GCN) kipf2016semi , a Graph Attention Network (GAT) that uses self-attention to weigh neighbor importance velivckovic2017graph , and a Graph Isomorphism Network (GIN), a highly expressive model effective at capturing complex structural motifs xu2018powerful .
5.2 Task 2: Spectrum-to-Property Prediction
The second, inverse task learns a mapping function $f:S\rightarrow P$, where $S$ is an XAS spectrum and $P$ is a set of local structural properties of the iron center. This demonstrates the dataset’s potential for extracting quantitative structural information directly from experimental spectra. The input $S$ is the 100-point interpolated XAS vector. The target properties $P$ are key structural descriptors: the coordination number (CN) of the iron atom (a classification task) and the mean nearest-neighbor distance (MNND) in Angstroms (a regression task).
We use two strong and interpretable baseline models for this task. The first is a standard Multi-Layer Perceptron (MLP), a feedforward neural network baseline where the final layer is adapted for either regression or classification. The second is a Random Forest (RF), a powerful tree-based ensemble method that is robust to overfitting and effective at capturing non-linear relationships in tabular data breiman2001random .
5.3 Experimental Setup
All models were trained and evaluated under a consistent framework. The dataset was randomly split into training (80%), validation (10%), and testing (10%) sets, with stratification applied to ensure a similar distribution of protein types and coordination numbers across sets. GNN models were implemented using the Deep Graph Library (DGL), while the MLP and RF models used PyTorch and Scikit-learn, respectively. Models were trained to minimize Mean Squared Error (MSE) for regression and Cross-Entropy Loss for classification, using the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$ and a scheduler to reduce the rate on a validation loss plateau. Training ran for a maximum of 300 epochs with an early stopping criterion. Performance was assessed on the held-out test set using Mean Absolute Error (MAE) for regression and Accuracy/Macro-F1 Score for classification.
5.4 Results
The trained baseline models provide a quantitative benchmark for the predictive utility of our dataset.
Structure-to-Spectrum Prediction
The performance of GNN models in predicting XAS spectra is summarized in Table 3. The Graph Isomorphism Network (GIN) demonstrated the best performance, achieving the lowest Mean Absolute Error (MAE) for both XANES and EXAFS prediction. This suggests its high expressive power is beneficial for capturing the complex relationship between the 3D atomic arrangement and the resulting spectral features. The Graph Attention Network (GAT) also performed competitively, outperforming the simpler GCN and indicating the advantage of learning to weigh the importance of different atomic neighbors.
Table 3: Performance of GNN models for Structure-to-Spectrum Prediction (MAE). Lower is better.
Model XANES MAE EXAFS MAE
GCN $0.085\pm 0.004$ $0.112\pm 0.005$
GAT $0.079\pm 0.003$ $0.105\pm 0.004$
GIN 0.075 $\pm$ 0.003 0.101 $\pm$ 0.004
Spectrum-to-Property Prediction
For the inverse task of predicting structural properties, the Random Forest (RF) model consistently outperformed the MLP baseline across all metrics (Table 4). For Coordination Number (CN) classification, the RF model achieved higher accuracy and F1-scores, suggesting its robustness in handling the spectral feature space. Similarly, for Mean Nearest-Neighbor Distance (MNND) regression, the RF yielded a lower MAE and a higher R2 value, indicating more reliable bond distance prediction. These results highlight the effectiveness of ensemble methods for tasks involving vector-based features like the interpolated spectra used here.
Table 4: Performance of classical models for Spectrum-to-Property Prediction.
Task Model Metric Value
CN Classification MLP Accuracy $72.4\pm 1.2\%$
F1-Score $70.1\pm 1.5\%$
Random Forest Accuracy 75.8 $\pm$ 1.1%
F1-Score 74.5 $\pm$ 1.3%
MNND Regression MLP MAE (Å) $0.065\pm 0.005$
R2 $0.91$
Random Forest MAE (Å) 0.058 $\pm$ 0.004
R2 0.93
Collectively, these baseline results validate the predictive potential of our dataset for both structure-to-spectrum and spectrum-to-property tasks, providing a strong quantitative foundation for the development of more sophisticated architectures.
6 Limitations
While our dataset and modeling framework offer a valuable resource for XAS-based learning, several limitations remain. First, the current study focuses exclusively on iron and Fe K-edge spectra, limiting generalizability across the periodic table. Expanding to other elements would broaden applicability for multi-element systems. Second, XAS spectra were manually digitized from published figures. Although expert validation was applied, minor inaccuracies may persist due to resolution limits or figure quality. Access to raw experimental data would enhance accuracy. Third, structural annotations were derived from diverse sources—PDB, CCDC, SMILES, and manual reconstruction—leading to variability in resolution and completeness. Some protein structures, especially those derived from textual or schematic descriptions, may be approximations rather than exact configurations. Fourth, the dataset is imbalanced, with fewer EXAFS samples relative to XANES, which may affect model performance on tasks requiring detailed oscillatory features. Lastly, our baseline models serve primarily as proof-of-concept. While effective, more advanced approaches—such as equivariant GNNs or physics-informed architectures—could yield further improvements and uncover deeper structure-spectrum relationships. Future efforts should address these limitations through dataset expansion, access to raw data, standardized structural extraction, and development of more sophisticated models.
7 Code and Data Availability
@@
The code used for processing the spectra and molecular structures is available at https://github.com/Airscker/XDIP. Including scripts for data interpolation, normalization, and basic spectral analysis, providing a starting point for researchers who are interested in further processing the data. The extracted dataset is available at https://airscker.github.io/XDIP. The software used for extracting numerical points from spectrum plots, i.e., WebPlotDigitizer V4, is open-source and available at https://github.com/automeris-io/WebPlotDigitizer.
8 Funding Disclosure
@@
We express our sincere thanks to Rutwik Segireddy and Keying Jia, for their work in assisting our dataset annotation. The work was supported in part by the Physical Biosciences Program and the Photochemistry and Biochemistry group within the US Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences (KC030402).
9 Competing interests
@@
Appendix A Technical Appendices and Supplementary Material
A.1 Detailed Data Collection and Curation
Literature Search and Selection
The data extraction pipeline is schematically overviewed in Figure 1 in the main paper. The first step involved constructing searching keywords to identify relevant literature. The search was conducted using combinations of keywords from two distinct sets: XAS, XANES, XAFS, EXAFS and Metalloprotein, Protein, Enzyme, Fe-iron. Each search term consisted of a pair of words, one from each set, such as "XAS Protein" or "EXAFS Metalloprotein".
These keywords were used to retrieve literature from various publishers, including Springer Nature, AAAS Science, American Chemical Society, Elsevier, Royal Society of Chemistry, Electrochemical Society, Wiley, MDPI, and others. To locate relevant literature, we utilized either the publishers’ search APIs or conducted manual searches under copyright permission. After excluding duplicate results, we had a pool of 20,915 articles (Table 5).
Table 5: Number of papers retrieved using different keyword combinations.
@@ A.1 Detailed Data Collection and Curation
From this pool, we manually selected papers that included both protein structure and the absorption spectra of iron elements. The selection criterion was the explicit mention of the protein and the inclusion of at least one Fe absorption spectrum. Eventually, we obtained 573 articles that met these criteria and were downloaded in PDF format for expert annotation.
Data Annotation and Refinement
After gathering the literature, human experts annotated the text and XAS plot information, converting it into digitized data samples. The extracted dataset was then refined by removing low-quality samples, poorly presented XAS and structures, and outdated annotations. This combination of automatic searching and manual data extraction and cleaning ensures the final dataset’s quality. Each data sample in our dataset consists of three main parts: (1) The entire or local protein structure (The first-coordination sphere of the element of interest), (2) The protein’s corresponding Fe K-edge XAS spectra, and (3) the basic information of papers from which the protein structure and XAS spectrum were derived.
A.2 XAS Extraction and Processing Details
Digitization from Published Plots
Extracting XAS data from the retrieved literature is a crucial step in constructing our database. XAS is divided into two regions: the near-edge spectra, also known as X-ray absorption near-edge structure (XANES), and the extended X-ray absorption fine structure (EXAFS), each providing specific information on the element studied. To extract numerical data points from the published XAS plots, we utilized the open-source data annotator WebPlotDigitizer V4. To guarantee the precision and reliability of data extraction, we followed the steps below:
In the process of extracting XAS data, we employ a meticulous digitization approach using the WebPlotDigitizer tool. Initially, we manually mask the spectral lines using the pencil tool to ensure precise alignment. The color recognition distance is set variably at 2, 5, or 10 pixels based on the clarity and overlap of the plotted lines; clearer plots are annotated at a denser resolution of 1 point per 2 pixels, while more complex, fuzzy plots require a broader setting of 1 point per 10 pixels. This flexibility allows us to capture data with high fidelity, respecting the original plot’s integrity without interpolation during the initial annotation phase.
Data Standardization and Metadata Annotation
Post-annotation, to standardize the spectral data for consistent analysis and comparison, we interpolate all spectra to a uniform length of 100 data points. This is based on the mean distribution of spectra lengths observed across the dataset, where both XANES and EXAFS lengths peak around 100, despite ranging up to 600. In particular, we also preserve our dataset’s original extracted data points. This dual provision enables researchers to select between the original or interpolated data depending on their specific analytical needs, thus offering flexibility while minimizing potential artifacts.
For the application of XAS in analysis, calibration energies are critical for accuracy. In our dataset, these energies are manually extracted from the text of each annotated paper. However, not all sources uniformly report this value; some mention using an iron foil standard without giving a specific energy, while others omit it entirely. In our efforts to maintain transparency, we have documented each instance of missing or incomplete information. Out of the dataset, 227 entries lack explicitly reported calibration energies. These cases are clearly labeled in our documentation files, enabling researchers to account for potential inconsistencies.
A.3 Detailed Data Record Format
Table 6: Format of each data record: description, key label, mandatory status, and data type.
@@ A.3 Detailed Data Record Format
The matrix of bond lengths between chemically bonded atoms Bond Lengths matrix of float numbers
The list of bond angles among neighboring bonds Bond Angles dictionary (key = angle vertex, value = float)
@@
The dataset developed in this study is hosted on https://airscker.github.io/XDIP. The format of the data records is detailed in Figure 2 in the main paper, which illustrates that each data sample consists of three parts: literature metadata, Fe K-edge XAS data, and the protein’s local structure containing Fe.
The literature metadata in each data record includes the Digital Object Identifier (DOI), the title of the research paper, and the XAS absorbing element, which is Fe in this work. These metadata serve as a quick reference for identifying the original research papers and their primary findings. Since most papers cover multiple sets of iron K-edge XAS and corresponding structures, we have stacked multiple absorption spectra and protein structures to capture all relevant details in the literature.
The protein structure includes atom types, their spatial coordinates, and their relationships. We extracted the protein’s structure or Fe-element neighborhood local structure through two methods: manual annotation of structure information and accessing data from public protein structure databases. The manually extracted protein (local) structures were derived through the following steps:
In contrast to manually extracted structures, extraction from open databases is more straightforward. In these cases, we simply include the URL for downloading the structure, the local path to the saved files, or the protein’s SMILES representation. Table 4 provides a detailed description of the data format, including every key label, whether it is mandatory, and its data type.
A.4 Documentation and Transparency Protocol
Table 7: Description of protein database documentation elements
@@ A.4 Documentation and Transparency Protocol
CCDC Cambridge Crystallographic Data Centre number, providing a reference to crystallographic data related to the protein.
@@
Comments Notes or comments about the quality or issues identified in the spectral or structural data, such as missing axes or unclear lines.
EXAFS Type Specifies the type of EXAFS data presented, whether it is raw signal data or Fourier transformed data.
@@
Another spreadsheet records the processes involved in extracting iron XAS spectra. It highlights various challenges encountered, such as missing axes on figures and ambiguities in figure interpretation. Such detailed records are crucial for enabling future researchers to understand and evaluate the decision-making processes affecting data inclusion or exclusion. Notably, we include and distinguish between EXAFS data represented in both k-space and r-space. This enhancement ensures that researchers can access and utilize the specific type of data they require for their analyses. We have meticulously labeled each dataset entry in our documentation files to reflect this categorization, addressing a critical need for clarity. The details of this documentation are explained in Table 8, and Figure 1(b) shows an example of an XAS spectrum data documentation log.
Data entries within our datasets were cross-verified against original publications to ensure accuracy and reliability. Spectral and structural data were, wherever feasible, compared against established databases to ensure consistency. Discrepancies and significant findings are carefully documented. The methodologies employed for data extraction are thoroughly detailed, including the tools used, such as the WebPlotDigitizer for converting plot images to numerical data. Each dataset entry is accompanied by extensive metadata such as detailed URLs of original publications, covering all aspects from data extraction to final dataset compilation. This approach not only enhances reproducibility but also facilitates the integration of our data with other datasets. It ensures a clear understanding of the context and specifics of the data collected.
(a) Example of protein structure data documentation
(b) Example of XAS spectrum data documentation
Figure S1: Examples of protein structure and XAS spectrum data documentation. (a) This example presents a data entry from our database, including the paper’s title, DOI, and URL. To accurately document the data, we included sections for both XANES and EXAFS to help identify the spectrum type corresponding to the protein structure. Due to missing descriptions in some papers, certain annotation elements, such as standards (which refer to the standard sample used) and ligation details (which provide information on ligands or binding groups associated with the protein structures), are omitted. Despite the missing information, we listed the samples mentioned in the paper that have both spectra and structures, with these key points shown in the sample section under the XANES and EXAFS parts. As outlined in the Data Records section, fields such as SMILES, PDB, and CCDC are selectively filled. (b) This table provides an example of an XAS spectrum data annotation log. We recorded the title of the literature, the location of the iron spectrum plot, its DOI, and URL. In addition to these key details, we included comments on the quality or issues identified in the spectral or structural data, such as missing axes or unclear lines. Furthermore, to ensure the accuracy and usability of our data, we highlighted the type of EXAFS spectrum extracted from the literature, noting whether the article presented Fourier-transformed EXAFS or not.